write an equation of the line that passes through (2,3) and perpendicular to 4x+5y=10
Answer by Alan3354(69443)

I just did one like that. Do it like this one. If you have trouble or want me to check your work email me at gsihoutx@aol.com
write an equation of the line in standard form that is perpendicular to 2x+6y=-4 and pass through (-1,-1)
-----------
1st, find the slope, m1, of 2x+6y = -4 by putting it in slope-intercept form. (That means solve for y)
2x+6y = -4
x+3y = -2
3y = -x - 2
y = (-1/3)x - 2
m1 = -1/3
---------
The slope of lines perpendicular is the negative inverse of m1, = +3
--------
Use y = mx + b and the point to find b.
-1 = 3*-1 + b
b = 2
--------
--> y = 3x + 2 is thru (-1,-1) and perpendicular to 2x+6y = -4